OpenStudy (anonymous):

find the polynomial with real coefficients of the smallest possible degree where i and 3 + i are zeros and the coefficient of the highest power is 1.

OpenStudy (anonymous):

ok well if the zeros are complex then they must come with their conjugates. zeros are therefore \[i, -i, 3+i,3-i\]

OpenStudy (anonymous):

in factored form you get \[(x-i)(x+i)(x-(3+i))(x-(3-i))\] and multiply out, which is not as bad as it looks

OpenStudy (anonymous):

first one is easy \[(x+i)(x-i)=x^2+1\]

OpenStudy (anonymous):

second one is not too bad since you always have \[(x-(a+bi))(x-(a-bi))=x^2-2ax+(a^2+b^2)\] so second one is \[x^2-6x+10\]

OpenStudy (anonymous):

and your last job is \[(x^2+1)(x^2-3x+10)\] which i will let you do

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

got it !

OpenStudy (anonymous):

good. i think i did the other one too. similar to this one

OpenStudy (anonymous):

okay , so ill have 3 seperate answers ? or just the last one..

OpenStudy (anonymous):

you have one answer but i did not do it. you still have to multiply out \[(x^2+1)(x^2-6x+10)\] "final answer" is \[x^4-6 x^3+11 x^2-6 x+10\]

OpenStudy (anonymous):

that is the polynomial you are looking for

OpenStudy (anonymous):

thats what i got , just checking. thanks a lot !